OpenStudy (anonymous):

solve the missing variable c = 2rrr, circumference of circle, c=7.42, r=3.14

OpenStudy (anonymous):

?

OpenStudy (anonymous):

elaborate your question

OpenStudy (anonymous):

@kid3

Directrix (directrix):

I'm:not clear on the question but I think it states: C = 2 * pi * r . Find if C=7.42 and pi =3.14 7.42 = 2* 3.14 * r 7.42 = 6.28 * r r = 7.42/6.28 Divide 7.42 by 6.28 to get r @kid3